OpenStudy (anonymous):

Find the equation in standard form which passes through (–1, 4) and (0, 9). You must show all work.

OpenStudy (rulnick):

slope = (9-4) / (0-(-1)) = 5 y-9 = 5(x-0) so 5x-y=-9

OpenStudy (anonymous):

(–1, 4) and (0, 9) m = 5 y - 9 = 5x => 5x -y = 9

OpenStudy (anonymous):

Oops, 5x -y = -9